OpenStudy (anonymous):

Please help solve! I have the first step but don't know what to do next!! 1. x^3-5x=0 I did x^3=5x first step 2. 2x^2-7x=15 I did 2x^2-7x-15=0 first step

OpenStudy (jdoe0001):

\(\bf x^3-5x=0 \implies x(x^2-5) = 0\)

OpenStudy (jdoe0001):

\(\bf 2x^2-7x=15 \implies 2x^2-7x-15 = 0\)

OpenStudy (anonymous):

x^3-5x=0 x(x^2-5)=0 either x=0 or x^2-5=0 or x^2=5 means x= sqrt(5) or -sqrt(5)

OpenStudy (jdoe0001):

next step on the second is to factor it to 2 binomials

OpenStudy (radar):

make that :